if the first term of a geometric sequence is 3/2 and the second and third terms are -3/4 and 3/8, respectively, which of the following represents the nth term of the sequence?
a) 3(-1)^n-1/2n
b) 3(-1)^n/2n
c) 3(-1)^n-1/2^n
d) 3(-1)^n/2^n
e) 3(-1)^n-1/2^n+1
please provide explanation!

Answer by stanbon(75887)

if the first term of a geometric sequence is 3/2 and the second and third terms are -3/4 and 3/8, respectively, which of the following represents the nth term of the sequence?
----
A geometric sequence has a common ratio between consecutive terms.
r = (-3/4)/(3/2) = -1/2
---
a(n) = ar^(n-1) = (3/2)(-1/2)^(n-1)
a(n) = 3*(-1)^(n-1)/2^(n)
================
Ans: c
==========
Cheers,
Stan H.
==========

a) 3(-1)^n-1/2n
b) 3(-1)^n/2n
c) 3(-1)^n-1/2^n
d) 3(-1)^n/2^n
e) 3(-1)^n-1/2^n+1